# Spherical maximal functions and Hardy spaces for Fourier integral operators

### v2

@@ Abstract.
We use the Hardy spaces for Fourier integral operators to obtain bounds for spherical maximal functions in $L^{p}(\mathbb{R}^{n})$, $n\geq 2$, where the radii of the spheres are restricted to a compact subset of $(0,\infty)$. These bounds extend to general hypersurfaces with non-vanishing Gaussian curvature, to the complex spherical means, and to geodesic spheres on compact manifolds. We also obtain improved maximal function bounds and pointwise convergence statements for wave equations, both on $\mathbb{R}^{n}$ and on compact manifolds. The maximal function bounds are essentially sharp for all $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty]$, for each such hypersurface, every complex spherical mean, and on every manifold.
@@
This research was funded in part by the National Science Center, Poland, grant 2021/43/D/ST1/00667. The third author is partially supported by NCN grant UMO-2023/49/B/ST1/01961.
@@ 1.1. Main results
Let $p\in[1,\infty]$ and $s>d(p)+\frac{1}{p}-\frac{n-1}{2}$. Then there exists a $C\geq 0$ such that
@@
for all $f\in C(\mathbb{R}^{n})\cap\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$.
In fact, (1.4) extends from $S^{n-1}$ to arbitrary compact hypersurfaces with non-vanishing Gaussian curvature, cf. Theorem 3.6, and a variable-coefficient version of Theorem 1.1 also holds; see Theorem 3.9. A consequence of the latter is a version of (1.4) for geodesic spheres on compact manifolds, Proposition 3.15. These results, and in particular Theorem 1.1, are essentially sharp for all $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty]$, as is shown in Theorems 4.8 and 4.9 and Proposition 4.12.
Another extension of Theorem 1.1 arises by considering maximal functions associated with the complex spherical means from [35]. In Proposition 3.11 we show that (1.4) holds for $s>d(p)+\frac{1}{p}-\frac{n-1}{2}-\operatorname{Re}(\alpha)$ if one replaces the average in (1.1) by the dilated complex spherical mean of order $\alpha\in\mathbb{C}$, with Theorem 1.1 corresponding to the case where $\alpha=0$. These bounds are essentially sharp for all $\alpha\in\mathbb{C}$ and $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty]$, as follows from Proposition 4.11.
@@
for all $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. Hence $e^{it\sqrt{-\Delta}}f(x)\to f(x)$ as $t\to 0$, for almost all $x\in\mathbb{R}^{n}$.
Theorem 1.2 is a special case of Corollary 3.2, and (1.5) is essentially sharp for all $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$, cf. Theorems 4.2 and 4.5. A version of Theorem 1.2 also holds on compact Riemannian manifolds $(M,g)$; see Proposition 3.14. In the latter case one replaces the Euclidean Laplacian $\Delta$ by the Laplace–Beltrami operator $\Delta_{g}$, and $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ by the space $\mathcal{H}^{s,p}_{FIO}(M)$ from Definition 2.5. On every such manifold, the resulting maximal function bounds are sharp for $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$, as is shown in Proposition 4.13.
@@ 1.2. Previous work
for $n\geq 3$ and $p>n/(n-1)$. It was later shown by Bourgain in [4] that (1.7) also holds for $n=2$ and $p>2$. Since (1.7) does not hold if $n=1$ and $p<\infty$, nor if $n\geq 2$ and $p\leq n/(n-1)$, this completely settled the question of the $L^{p}$ boundedness of $\mathcal{M}_{(0,\infty)}$. Indeed, apart from the fact that the condition $p>n/(n-1)$ is sharp, for such $p$ one cannot improve (1.7) by considering a larger space than $L^{p}(\mathbb{R}^{n})$ on the right-hand side of (1.7). This follows from the fact that the spherical averages converge pointwise almost everywhere to a multiple of $f$ as $t\to 0$.
@@
for $n\geq 3$, $p>n/(n-1)$ and suitable $s<0$, follow from the work of Sogge and Stein in [34]. In fact, they obtained a version of (1.7) for more general hypersurfaces in $\mathbb{R}^{n}$, by combining bounds as in (1.8) with a scaling argument. A similar approach was used in the work [23] of Mockenhoupt, Seeger and Sogge, who relied on improved bounds for $\|\mathcal{M}_{[1,2]}f\|_{L^{p}(\mathbb{R}^{2})}$ to give a simplied proof of (1.7) for $n=2$ and $p>2$ (see also [32, 24] for more general hypersurfaces in $\mathbb{R}^{2}$). Estimates as in (1.8) are also strongly related to the frequently studied question of the $L^{p}$ boundedness of maximal functions associated with the complex spherical means (see e.g. [23, 21, 22]). Finally, in [28] Schlag obtained sharp $L^{p}\mapsto L^{q}$ bounds for $\mathcal{M}_{[1,2]}$, with $q>p$, which are related to (1.8) due to Sobolev embeddings (see also [29, 18]). However, for general $p$ it does not appear to be known what the optimal condition on $s$ is for (1.8) to hold.
@@
are known to hold for $p\geq 2(n+1)/(n-1)$ and $s>2s(p)$, as a corollary of [6] (see [1, 2] for the analogous statement on compact manifolds). For $n=2$ one can obtain (1.9) for $p>2n/(n-1)$ and $s>2s(p)$ by relying on [12], and similarly for compact Riemannian surfaces, using [11]. These bounds are all essentially sharp in terms of the exponent $s$. Moreover, in [15] the endpoint case of (1.9) was obtained for $n\geq 4$, $p>2+\frac{4}{n-3}>2\frac{n+1}{n-1}$ and $s=2s(p)$, with an analogous statement on compact manifolds being contained in [19]. Nonetheless, for general $p$ it does not appear to be known what the optimal condition on $s$ is for (1.9) to hold.
@@ 1.3. Discussion of the results
In particular, for $1<p<\infty$, Theorem 1.1 yields (1.8) with $s>d(p)+s(p)+\frac{1}{p}-\frac{n-1}{2}$, and Theorem 1.2 yields (1.9) with $s>d(p)+s(p)+\frac{1}{p}$. One has $\mathcal{H}^{2}_{FIO}(\mathbb{R}^{n})=L^{2}(\mathbb{R}^{n})$, so for $p=2$ all our results apply to $W^{s,2}(\mathbb{R}^{n})$. However, for $p\neq 2$ the corollaries of Theorems 1.1 and 1.2 for $W^{s,p}(\mathbb{R}^{n})$ are strictly weaker than the estimates in Theorems 1.1 and 1.2 themselves, since $W^{s(p),p}(\mathbb{R}^{n})$ is a proper subspace of $\mathcal{H}^{p}_{FIO}(\mathbb{R}^{n})$. In fact, the Sobolev exponents in (1.10) are sharp, and therefore the bounds in Theorems 1.1 and 1.2 are significantly stronger than their corollaries regarding (1.8) and (1.9), at least for certain initial data. This is particularly relevant for $p\geq 2(n+1)/(n-1)$, where one has $d(p)+s(p)+\frac{1}{p}=2s(p)$ and the resulting condition on $s$ in (1.8) and (1.9) is essentially sharp, due to the mapping properties of the individual spherical means and half-wave propagators. Theorems 1.1 and 1.2 show that (1.8) and (1.9) can nonetheless be improved, by working with a different space of initial data. We also note that, for suitable Fourier integral operators $T$, one can replace $f$ by $Tf$ on the left-hand sides of (1.4) and (1.5), whereas this is not possible when working directly with (1.8) and (1.9).
For $n=2$ and $p>2n/(n-1)$, one can use [12] to obtain (1.8) and (1.9) for $s>2s(p)-\frac{n-1}{2}$ and $s>2s(p)$, respectively. Due to the sharpness of the Sobolev exponents in (1.10), for $2n/(n-1)<p<2(n+1)/(n-1)$ these bounds do not follow from Theorems 1.1 and 1.2, nor does the converse implication hold. Also, the endpoint version of (1.9) in [15] neither follows from Theorem 1.2 , nor the other way around. A similar comparison holds for the analogues of Theorems 1.1 and 1.2 on compact manifolds, Propositions 3.15 and 3.14, and the results in [11, 19].
@@ 1.4. Ideas behind the proofs
A significant part of the article is dedicated to proving that our results are essentially sharp. In the case of Theorem 1.2, sharpness follows for $p\geq 2(n+1)/(n-1)$ by combining (1.10) with the $L^{p}$ mapping properties of the individual operators, but in Theorem 4.2 we prove a more general lower bound for $s$ in (1.5) which does not hold in the $L^{p}$ scale (see Remark 4.4). To prove sharpness for $1\leq p\leq 2$ one has to produce additional focusing of order $1/p$, which we do in Theorem 4.5 using a Knapp example that cannot be used to prove sharpness of the corresponding estimate in the $L^{p}$ scale (see Remark 4.7). In fact, Theorems 4.2 and 4.5 both concern more general Fourier integral operators, and the proofs rely on an improvement of a lemma from [27, 20] which relates such operators to associated flow maps, Lemma 4.1. We do not know whether our results are sharp for $2<p<2(n+1)/(n-1)$ (see also Remark 4.6).
@@
Finally, given that the assumptions for the sharpness statements in Theorems 4.2 and 4.5 are of a local nature, it suffices to work in local coordinates when deriving sharpness of our results on manifolds. There the required statements follows from the sharpness of our other results.
@@ 1.6. Notation and terminology
The natural numbers are =$\mathbb{N}=\{1,2,\ldots\}$, and +:=$\mathbb{Z}_{+}:=\mathbb{N}\cup\{0\}$. Throughout, we fix an $n\in\mathbb{N}$ with $n\geq 2$.
@@ 2.2. Results on $\mathbb{R}^{n}$
for all $p\in[1,\infty]$ and $s\in\mathbb{R}$, as follows from [13, Theorem 7.4]. Here $s(p)$ is as in (1.2). If $p<\infty$, then the Schwartz functions lie dense in $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ (see [13, Proposition 6.6]). Although we will not make explicit use of the following properties in the present article, we also note that the Hardy spaces for FIOs form a complex interpolation scale, cf. [13, Proposition 6.7], and that $(\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n}))^{*}=\mathcal{H}^{-s,p^{\prime}}_{FIO}(\mathbb{R}^{n})$ for all $p\in[1,\infty)$ and $s\in\mathbb{R}$, cf. [13, Proposition 6.8]. By [31, 13, 20], $\mathcal{H}^{s,1}_{FIO}(\mathbb{R}^{n})$ has suitable molecular and atomic decompositions.
The following proposition is a slight generalization of [27, Corollary 3.6 and Theorem 4.4], to the case where the phase function is not defined globally and the symbol is time dependent. Recall that a subset $V\subseteq\mathbb{R}^{n}\setminus\{0\}$ is conic if $\lambda\eta\in V$ for all $\eta\in V$ and $\lambda>0$, and recall the definition of the exponent $d(p)$ from (1.3).
@@
Let $\chi\in C^{\infty}(\mathbb{R}^{n}\setminus\{0\})$ be real-valued and positively homogeneous of degree zero, and such that $\mathrm{supp}(\chi)\subseteq V$ and $\chi\equiv 1$ on $V^{\prime}$. Then $\phi^{\prime}:=\phi\chi\in C^{\infty}(\mathbb{R}^{n}\setminus\{0\})$ is real-valued and positively homogeneous of degree $1$. Moreover, $\mathrm{supp}(a(t\cdot))\subseteq V^{\prime}\cup\{0\}$ for all $t>0$, since $V^{\prime}$ is conic, and $\{a(t\cdot)\mid t\in I\}$ is a uniformly bounded collection in $S^{m}(\mathbb{R}^{n})$. Hence (1) follows by applying [27, Lemma 3.5] to $e^{it\phi^{\prime}(D)}a(tD)\langle D\rangle^{-m}=e^{it\phi(D)}a(tD)\langle D\rangle^{-m}$, for $t\in I$.
On the other hand, the proof of (2) is analogous to that of [27, Theorem 4.4], relying in particular on the $\ell^{p}$ decoupling inequality from [7], for hypersurfaces with non-vanishing Gaussian curvature. Note that the condition that $\mathrm{rank}\,\partial_{\eta\eta}^{2}\phi(\eta)=n-1$ for all $\eta\in V$ is equivalent to $\{\eta\in V\mid\phi(\eta)=1\}$ having non-vanishing Gaussian curvature. ∎
@@
The result is contained in [20]. More precisely, (1) is [20, Corollary 2.4 and Remark 2.6], and (2) can be deduced from [20, Corollary 5.4] in the same way as in the proof of [20, Theorem 5.7]. ∎
@@
Recall that, for $(z_{0},\eta_{0})\in\mathrm{supp}(a)$ with $\eta_{0}\neq 0$, the generalized Gauss map is $G(z_{0},\eta_{0})=G_{0}(z_{0},\eta_{0})/|G_{0}(z_{0},\eta_{0})|$, where $G_{0}(z_{0},\eta_{0}):=\wedge_{i=1}^{n}\partial_{\eta_{i}}\partial_{z}\Phi(z_{0},\eta_{0})$. In other words, (2.7) says that the unit normal vectors ±$\pm\theta$ to the conical hypersurface $\{\partial_{z}\Phi(z_{0},\lambda\eta)\mid(z_{0},\eta)\in\mathrm{supp}(a),\lambda>0\}\subseteq\mathbb{R}^{n+1}$ at $\partial_{z}\Phi(z_{0},\eta_{0})$ satisfy
@@
This is in turn equivalent to $\{\partial_{z}\Phi(z_{0},\lambda\eta)\mid(z_{0},\eta)\in\mathrm{supp}(a),\lambda>0\}\subseteq\mathbb{R}^{n+1}$ having $n-1$ non-vanishing principal curvatures, which, in the terminology of [32], means that the canonical relation of $T$ satisfies the cinematic curvature condition.
Note also that, if (2.7) holds, then one cannot have $\partial_{t}\Phi(x_{0},t,\eta_{0})|_{t=t_{0}}\neq 0$ for all $(z_{0},\eta_{0})$ in an open subset of $\mathrm{supp}(a)$, where we write $z_{0}=(x_{0},t_{0})$. Indeed, otherwise an open subset of the conical hypersurface $\{\partial_{z}\Phi(z_{0},\lambda\eta)\mid(z_{0},\eta)\in\mathrm{supp}(a),\lambda>0\}$ would be contained in the plane where the last coordinate vanishes, which would prohibit it from having $n-1$ nonvanishing principal curvatures.
We will also use frequently that $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ is invariant under changes of coordinates, for all $1\leq p\leq\infty$ and $s\in\mathbb{R}$, as follows from [13, Theorem 6.10] (see also [27, Proposition 3.3]).

### v3

@@ Abstract.
We use the Hardy spaces for Fourier integral operators to obtain bounds for spherical maximal functions in $L^{p}(\mathbb{R}^{n})$, $n\geq 2$, where the radii of the spheres are restricted to a compact interval in $(0,\infty)$. These bounds extend to general hypersurfaces with non-vanishing Gaussian curvature, to the complex spherical means, and to geodesic spheres on compact manifolds. We also obtain improved maximal function bounds and pointwise convergence statements for wave equations, both on $\mathbb{R}^{n}$ and on compact manifolds. The maximal function bounds are essentially sharp for all $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$, for each such hypersurface, every complex spherical mean, and on every manifold.
@@
This research was funded in part by the National Science Center, Poland, grant 2021/43/D/ST1/00667. N. J. Liu is supported by China Postdoctoral Science Foundation (No. 2024M763732) and NNSF of China (No. 12501132). J. Rozendaal is partially supported by NCN grant UMO-2023/49/B/ST1/01961. L. Song is supported by NNSF of China (No. 12471097).
@@ 1.1. Main results
Let $p\in[1,\infty)$ and $s>d(p)+\frac{1}{p}-\frac{n-1}{2}$. Then there exists a $C\geq 0$ such that
@@
for all $f\in C(\mathbb{R}^{n})\cap\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. Conversely, if $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$ and (1.4) holds for all $f\in\mathcal{S}(\mathbb{R}^{n})$, then $s\geq d(p)+\frac{1}{p}-\frac{n-1}{2}$.
In fact, (1.4) extends from $S^{n-1}$ to arbitrary compact hypersurfaces with non-vanishing Gaussian curvature, cf. Theorem 3.7, and a variable-coefficient version also holds; see Theorem 3.10. A consequence of the latter is an extension of (1.4) to geodesic spheres on compact manifolds, Proposition 3.16. These results are all essentially sharp for $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$, as is shown in Theorems 4.8 and 4.9 and Proposition 4.12. See also Remark 3.4 for the case where $p=\infty$.
Another extension of Theorem 1.1 arises by considering maximal functions associated with the complex spherical means from [39]. In Proposition 3.12 we show that (1.4) holds for $s>d(p)+\frac{1}{p}-\frac{n-1}{2}-\operatorname{Re}(\alpha)$ if one replaces the average in (1.1) by the dilated complex spherical mean of order $\alpha\in\mathbb{C}$, with Theorem 1.1 corresponding to the case where $\alpha=0$. These bounds are essentially sharp for all $\alpha\in\mathbb{C}$ and $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$, as follows from Proposition 4.11.
@@
for all $f\in\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$. Hence $e^{it\sqrt{-\Delta}}f(x)\to f(x)$ as $t\to 0$, for almost all $x\in\mathbb{R}^{n}$. Conversely, if $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$ and (1.5) holds for all $f\in\mathcal{S}(\mathbb{R}^{n})$, then $s\geq d(p)+\frac{1}{p}$.
Theorem 1.2 is a special case of Corollary 3.3 and Theorems 4.2 and 4.5, with the latter two results containing the sharpness statement. A version of Theorem 1.2 also holds on compact Riemannian manifolds $(M,g)$; see Proposition 3.15. In this case one replaces the Euclidean Laplacian $\Delta$ by the Laplace–Beltrami operator $\Delta_{g}$, and $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ by the space $\mathcal{H}^{s,p}_{FIO}(M)$ from Definition 2.5. On every such manifold, the resulting maximal function bounds are essentially sharp for $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$, as is shown in Proposition 4.13.
@@ 1.2. Previous work
for $n\geq 3$ and $p>\frac{n}{n-1}$. It was later shown by Bourgain in [4] that (1.7) also holds for $n=2$ and $p>2$. Since (1.7) does not hold if $n=1$ and $p<\infty$, nor if $n\geq 2$ and $p\leq\frac{n}{n-1}$, this completely settled the question of the $L^{p}$ boundedness of $\mathcal{M}_{(0,\infty)}$. Indeed, apart from the fact that the condition $p>\frac{n}{n-1}$ is sharp, for such $p$ one cannot improve (1.7) by considering a larger space than $L^{p}(\mathbb{R}^{n})$ on the right-hand side of (1.7). This follows from the fact that the spherical averages converge pointwise almost everywhere to a multiple of $f$ as $t\to 0$.
@@
for $n\geq 3$, $p>\frac{n}{n-1}$ and suitable $s<0$, follow from the work of Sogge and Stein in [38]. In fact, they obtained a version of (1.7) for more general hypersurfaces in $\mathbb{R}^{n}$, by combining bounds as in (1.8) with a scaling argument, and their techniques lead to the condition $s>\frac{1}{p}-(n-1)\min(\frac{1}{p},\frac{1}{p^{\prime}})$ in (1.8) (see [38, pp. 182-183]). A similar approach was used in the work [25] of Mockenhoupt, Seeger and Sogge, who relied on improved bounds for $\|\mathcal{M}_{[1,2]}f\|_{L^{p}(\mathbb{R}^{2})}$ to give a simplified proof of (1.7) for $n=2$ and $p>2$ (see also [36, 26] for more general hypersurfaces in $\mathbb{R}^{2}$). Schlag [32] obtained sharp $L^{p}\mapsto L^{q}$ bounds for $\mathcal{M}_{[1,2]}$, with $q>p$. These are stronger than $L^{p}\to L^{p}$ bounds and are related to (1.8) due to Sobolev embeddings (see [33, 21, 20] as well). We also refer to e.g. [29] for results on spherical maximal functions in the case where the radii are restricted to a fractal subset of $[1,2]$. Finally, estimates as in (1.8) are connected to the frequently studied question of the $L^{p}$ boundedness of maximal functions associated with the complex spherical means (see e.g. [25, 23, 24]). However, it is crucial for the present article to note that, for general $p$, it does not appear to be known what the optimal condition on $s$ is for (1.8) to hold.
@@
are known to hold for $p\geq\frac{2(n+1)}{n-1}$ and $s>2s(p)$, as a corollary of [6] (see [1, 2] for the analogous statement on compact manifolds). For $n=2$ one can obtain (1.9) for $p>\frac{2n}{n-1}$ and $s>2s(p)$ by relying on [15], and similarly for compact Riemannian surfaces, using [14]. These bounds are all essentially sharp in terms of the exponent $s$. For $p=2$ and $s>1/2$, (1.9) is due to Cowling [9], and here the condition $s>1/2$ is necessary [42]. In particular, given that $\mathcal{H}^{s,2}_{FIO}(\mathbb{R}^{n})=W^{s,2}(\mathbb{R}^{n})$, this corresponds to the case $p=2$ in Theorem 1.2. We also refer to [28, 8] for endpoint bounds in the setting where $p$ equals $2$ on the right-hand side, but not on the left-hand side, of (1.9). However, for general $p$ it does not appear to be known what the optimal condition on $s$ is for (1.9) to hold.
@@ 1.3. Discussion of the results
In particular, for $1<p<\infty$, Theorem 1.1 yields (1.8) with $s>d(p)+s(p)+\frac{1}{p}-\frac{n-1}{2}$, and Theorem 1.2 yields (1.9) with $s>d(p)+s(p)+\frac{1}{p}$. One has $\mathcal{H}^{2}_{FIO}(\mathbb{R}^{n})=L^{2}(\mathbb{R}^{n})$, so for $p=2$ all our results apply to $W^{s,2}(\mathbb{R}^{n})$. However, for $p\neq 2$ the corollaries of Theorems 1.1 and 1.2 for $W^{s,p}(\mathbb{R}^{n})$ are strictly weaker than the estimates in Theorems 1.1 and 1.2 themselves, since $W^{s(p),p}(\mathbb{R}^{n})$ is a proper subspace of $\mathcal{H}^{p}_{FIO}(\mathbb{R}^{n})$. In fact, the Sobolev exponents in (1.10) are sharp, and therefore the bounds in Theorems 1.1 and 1.2 are significantly stronger than their corollaries regarding (1.8) and (1.9), at least for certain initial data. This is particularly relevant for $p\geq\frac{2(n+1)}{n-1}$, where one has $d(p)+s(p)+\frac{1}{p}=2s(p)$ and the resulting condition on $s$ in (1.8) and (1.9) is essentially sharp, due to the mapping properties of the individual spherical means and half-wave propagators. Theorems 1.1 and 1.2 show that (1.8) and (1.9) can nonetheless be improved, by working with a different space of initial data. We also note that, for suitable Fourier integral operators $T$, one can replace $f$ by $Tf$ on the left-hand sides of (1.4) and (1.5), whereas this is not possible when working directly with (1.8) and (1.9).
For $n=2$ and $p>\frac{2n}{n-1}$, one can use [15] to obtain (1.8) and (1.9) for $s>2s(p)-\frac{n-1}{2}$ and $s>2s(p)$, respectively. Due to the sharpness of the Sobolev exponents in (1.10), for $\frac{2n}{n-1}<p<\frac{2(n+1)}{n-1}$ these bounds do not follow from Theorems 1.1 and 1.2, nor does the converse implication hold. A similar comparison holds for the analogues of Theorems 1.1 and 1.2 on compact manifolds, Propositions 3.16 and 3.15, and the results in [14].
While the sharp $L^{p}\to L^{q}$ improving properties from [32, 33, 21] only hold for $\frac{n}{n-1}<p\leq q$, Theorems 1.1 and 1.2 yield sharp bounds for all $p\in[1,2]\cup[\frac{2(n+1)}{n-1},\infty)$. In fact, some of our estimates involving $\mathcal{H}^{p}_{FIO}(\mathbb{R}^{n})$ are sharp even when the analogous bounds for $L^{p}(\mathbb{R}^{n})$ can be improved (see e.g. Remark 4.4).
@@ 1.4. Ideas behind the proofs
Although our approach to obtaining maximal function bounds is analogous to that in the literature, one of the main goals of this article is to prove estimates which are sharp in a variety of geometric settings, and a significant part of the article is dedicated to showing that our results are indeed essentially sharp. As in the case of the results on local smoothing in [31, 22], working with the Hardy spaces for Fourier integral operators allows one to obtain sharp estimates even where the corresponding sharp bounds for $L^{p}(\mathbb{R}^{n})$ are not yet known, at the price of the additional work required to deal with the more complicated $\mathcal{H}^{p}_{FIO}(\mathbb{R}^{n})$ norm.
In the case of Theorem 1.2, sharpness follows for $p\geq\frac{2(n+1)}{n-1}$ by combining (1.10) with the $L^{p}$ mapping properties of the individual operators, but in Theorem 4.2 we prove a more general lower bound for $s$ in (1.5) which does not hold in the $L^{p}$ scale (see Remark 4.4). To prove sharpness for $1\leq p\leq 2$ one has to produce additional focusing of order $1/p$, which we do in Theorem 4.5 using a Knapp example that cannot be used to prove sharpness of the corresponding estimate in the $L^{p}$ scale (see Remark 4.7). In fact, Theorems 4.2 and 4.5 both concern more general Fourier integral operators, and the proofs rely on an improvement of a lemma from [31, 22] which relates such operators to associated flow maps, Lemma 4.1. We do not know whether our results are sharp for $2<p<\frac{2(n+1)}{n-1}$ (see also Remark 4.6).
@@
Finally, given that the assumptions for the sharpness statements in Theorems 4.2 and 4.5 are of a local nature, it suffices to work in local coordinates when deriving sharpness of our results on manifolds. There one can use that our other results are sharp under very mild conditions.
@@ 1.6. Notation and terminology
The natural numbers are $\mathbb{N}=\{1,2,\ldots\}$, and $\mathbb{Z}_{+}:=\mathbb{N}\cup\{0\}$. Throughout, we fix an $n\in\mathbb{N}$ with $n\geq 2$.
@@ 2.2. Results on $\mathbb{R}^{n}$
for all $p\in[1,\infty]$ and $s\in\mathbb{R}$, as follows from [16, Theorem 7.4]. Here $s(p)$ is as in (1.2). If $p<\infty$, then the Schwartz functions lie dense in $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ (see [16, Proposition 6.6]). Although we will not make explicit use of the following properties in the present article, we also note that the Hardy spaces for FIOs form a complex interpolation scale, cf. [16, Proposition 6.7], and that $(\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n}))^{*}=\mathcal{H}^{-s,p^{\prime}}_{FIO}(\mathbb{R}^{n})$ for all $p\in[1,\infty)$ and $s\in\mathbb{R}$, cf. [16, Proposition 6.8]. By [35, 16, 22], $\mathcal{H}^{s,1}_{FIO}(\mathbb{R}^{n})$ has suitable molecular and atomic decompositions.
The following proposition is a slight generalization of [31, Corollary 3.6 and Theorem 4.4], to the case where the phase function is not defined globally and the symbol is time dependent. Recall that a subset $V\subseteq\mathbb{R}^{n}\setminus\{0\}$ is conic if $\lambda\eta\in V$ for all $\eta\in V$ and $\lambda>0$, and recall the definition of the exponent $d(p)$ from (1.3).
@@
Let $\chi\in C^{\infty}(\mathbb{R}^{n}\setminus\{0\})$ be real-valued and positively homogeneous of degree zero, and such that $\mathrm{supp}(\chi)\subseteq V$ and $\chi\equiv 1$ on $V^{\prime}$. Then $\phi^{\prime}:=\phi\chi\in C^{\infty}(\mathbb{R}^{n}\setminus\{0\})$ is real-valued and positively homogeneous of degree $1$. Moreover, $\mathrm{supp}(a(t\cdot))\subseteq V^{\prime}\cup\{0\}$ for all $t>0$, since $V^{\prime}$ is conic, and $\{a(t\cdot)\mid t\in I\}$ is a uniformly bounded collection in $S^{m}(\mathbb{R}^{n})$. Hence (1) follows by applying [31, Lemma 3.5] to $e^{it\phi^{\prime}(D)}a(tD)\langle D\rangle^{-m}=e^{it\phi(D)}a(tD)\langle D\rangle^{-m}$, for $t\in I$.
On the other hand, the proof of (2) is analogous to that of [31, Theorem 4.4], relying in particular on the $\ell^{p}$ decoupling inequality from [7], for hypersurfaces with non-vanishing Gaussian curvature. Note that the condition that $\mathrm{rank}\,\partial_{\eta\eta}^{2}\phi(\eta)=n-1$ for all $\eta\in V$ is equivalent to $\{\eta\in V\mid\phi(\eta)=1\}$ having non-vanishing Gaussian curvature. ∎
@@
The result is contained in [22]. More precisely, (1) is [22, Corollary 2.4 and Remark 2.6], and (2) can be deduced from [22, Corollary 5.4] in the same way as in the proof of [22, Theorem 5.7]. ∎
@@
Recall that, for $(z_{0},\eta_{0})\in\mathrm{supp}(a)$ with $\eta_{0}\neq 0$, the generalized Gauss map is $G(z_{0},\eta_{0})=G_{0}(z_{0},\eta_{0})/|G_{0}(z_{0},\eta_{0})|$, where $G_{0}(z_{0},\eta_{0}):=\wedge_{i=1}^{n}\partial_{\eta_{i}}\partial_{z}\Phi(z_{0},\eta_{0})$. In other words, (2.7) says that the unit normal vectors $\pm\theta$ to the conical hypersurface $\{\partial_{z}\Phi(z_{0},\lambda\eta)\mid(z_{0},\eta)\in\mathrm{supp}(a),\lambda>0\}\subseteq\mathbb{R}^{n+1}$ at $\partial_{z}\Phi(z_{0},\eta_{0})$ satisfy
@@
This is in turn equivalent to $\{\partial_{z}\Phi(z_{0},\lambda\eta)\mid(z_{0},\eta)\in\mathrm{supp}(a),\lambda>0\}\subseteq\mathbb{R}^{n+1}$ having n1n-1 non-vanishing principal curvatures, which, in the terminology of [36], means that the canonical relation of $T$ satisfies the cinematic curvature condition.
Note also that, if (2.7) holds, then one cannot have $\partial_{t}\Phi(x_{0},t,\eta_{0})|_{t=t_{0}}=0$ for all $(z_{0},\eta_{0})$ in an open subset of $\mathrm{supp}(a)$, where we write $z_{0}=(x_{0},t_{0})$. Indeed, otherwise an open subset of the conical hypersurface $\{\partial_{z}\Phi(z_{0},\lambda\eta)\mid(z_{0},\eta)\in\mathrm{supp}(a),\lambda>0\}$ would be contained in the plane where the last coordinate vanishes, which would prohibit it from having n1n-1 nonvanishing principal curvatures.
We will also use frequently that $\mathcal{H}^{s,p}_{FIO}(\mathbb{R}^{n})$ is invariant under changes of coordinates, for all $1\leq p\leq\infty$ and $s\in\mathbb{R}$, as follows from [16, Theorem 6.10] (see also [31, Proposition 3.3]).